OpenStudy (anonymous):

2x + 12 = 18 - x X = A 3 B 10 C 6 D 2 8x - 3 = 15x + 18 X = A -3 B 3 C 2.1 D .9 6y - 6 = 4y + 16 Y = A 2.2 B -2.2 C 11 D 5 3(x - 4) = 5(x + 2) X = A 11 B -11 C 1 D -1 Medal Given for best help :)

OpenStudy (anonymous):

2x + 12 = 18 - x (add x to both sides) 2x + x + 12 = 18 (subtract 12 from both sides) 2x + x = 18 - 12 (combine like terms) 3x = 18 (divide both sides by 3) x = 6 =============== 8x - 3 = 15x + 18 (subtract 15x from both sides) 8x - 3 - 15x = 18 (add 3 to both sides) 8x - 15x = 18 + 3 (combine like terms) -7x = 21 (divide both sides by -7) x = -3 =============== 6y - 6 = 4y + 16 (subtract 4y from both sides) 6y - 6 - 4y = 16 (add 6 to both sides) 6y - 4y = 16 + 6 (combine like terms) 2y = 22 (divide both sides by 2) y = 11 =============== 3(x - 4) = 5(x + 2) (distribute through the parenthesis) 3x - 12 = 5x + 10 (subtract 5x from both sides) 3x - 12 - 5x = 10 (add 12 to both sides) 3x - 5x = 10 + 12 (combine like terms) -2x = 22 (divide by -2) x = -11

OpenStudy (anonymous):

THANK YOU!!!!

OpenStudy (anonymous):

The first one was actually 2 though (that's what it said when I checked it). Thanks for the other ones!